the birth weight of babies is normally distributed with mean 3500g and standard deviation 500g. what is the probability that a baby is born that weighs less than 3100g? construct the standard normal distribution density curve, then shade your seeking area
Answer #1

Answer:

mean = 3500

standard deviation = 500

P(X < 3100) = P((x-u)/s < (3100 - 3500)/500)

= P(z < -400/500)

= P(z < - 0.8)

= 0.2118554 [since from z table]

= 0.2119